OpenStudy (anonymous):

What are the roots of the equation 3x^2+5x=8

OpenStudy (anonymous):

Subtract both sides by 8, then use quadratic formula.

OpenStudy (anonymous):

Whats the formula? What i did was subtract then factored to (2x+5)(x+1)

OpenStudy (anonymous):

Ah, or you can factor it. You did it, now find the x. You know how to?

OpenStudy (anonymous):

I think i have but i dont know if its correct i got x > -2.5 && x > - 1

OpenStudy (anonymous):

it is (-8/3) and 1

OpenStudy (anonymous):

i used my graphing calculator application to find it

OpenStudy (anonymous):

I dont just want the answer i want to know how to do it

OpenStudy (anonymous):

use the quadratic formular

OpenStudy (anonymous):

what does it mean? @micahwood50

OpenStudy (anonymous):

Don't listen to him, (2x+5)(x+1) = 0 so either (2x+5) or (x+1) is equal to zero so x can be -1 for (x+1) to be zero and x also can be -5/2 or -2.5

hartnn (hartnn):

@micahwood50 factoring was not done correctly. if aske needs quadratic formula, then here it is: Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{{b}^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

then it is 3x^2 +5x -8=0 a=3 b=5 c=-8

OpenStudy (anonymous):

now substitute the values in

OpenStudy (anonymous):

\[\frac{ -5 \pm \sqrt{{5}^2-4(3)(-8)} }{2(3) }\]

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

now calculate \[\frac{ -5\pm \sqrt{25-4(-24)} }{ 6 }\]

OpenStudy (anonymous):

you got the idea now

OpenStudy (anonymous):

i plugged in and it doesnt work

OpenStudy (radar):

"Whats the formula? What i did was subtract then factored to (2x+5)(x+1)" Would of worked if you had factored correctly. Factors to (x-1)(3x+8)=0 x=1 x=-8/3

OpenStudy (radar):

Factoring if possible, is simpler than quadratic formula, but, factoring is not always simple!

OpenStudy (radar):

........and not always possible lol

OpenStudy (radar):

@leanswag , the question is still open, are you still having problems with it?